To what temperature would you have to heat a brass rod for it to be 1.7%longer than it is at 20 ?C?
Answer #1

linear thermal expansion coef brass 19e-6 /K
?L/L = ??T, ? is linear thermal expansion coef
L = L?(1 + ??T)

?L/L = ??T
??T = 0.017
?T = 0.017 / 19e-6 = 895
Answer #2


linear thermal expansion coef brass 19e-6 /K

?L/L = ??T, ? is linear thermal expansion coef
1.7% is 0.017
0.017 = 19*10^-6 * ?T
?T = 894.73
so T = 894.73 +20 = 914.73 C

You can get different numbers, as the coef will vary with different types of brass.